# Self Model for Embodied Intelligence: Modeling Full-Body Human Musculoskeletal System and Locomotion Control with Hierarchical Low-Dimensional Representation

### v2

@@ Abstract
Modeling and control of the human musculoskeletal system is important for understanding human motor functions, developing embodied intelligence, and optimizing human-robot interaction systems. However, current open-source models are restricted to a limited range of body parts and often with a reduced number of muscles. There is also a lack of algorithms capable of controlling over 600 muscles to generate reasonable human movements. To fill this gap, we build a musculoskeletal model with 90 body segments, 206 joints, and 700 muscle-tendon units, allowing simulation of full-body dynamics and interaction with various devices. We develop a new algorithm using low-dimensional representation and hierarchical deep reinforcement learning to achieve state-of-the-art full-body control. We validate the effectiveness of our model and algorithm in simulations with real human locomotion data. The musculoskeletal model, along with its control algorithm, will be made available to the research community to promote a deeper understanding of human motion control and better design of interactive robots.
@@ I Introduction
Many of our daily movements require the use of the entire body. For instance, walking and running involve the coordinated movement of the lower limbs for propulsion, complemented by the swinging motion of the upper limbs for balance, while the trunk muscles contribute to torso stability. However, there is no open-source musculoskeletal model that represents the entire human body. Constructing a full-body human musculoskeletal model enables comprehensive characterization of the action space and state space in human movement simulations. The dynamics of the model improves our understanding of human motor intelligence. Furthermore, current algorithms struggle to address the highly dimensional control problem associated with full-body human musculoskeletal motion.
@@
Figure 1: (a) Our full-body musculoskeletal model. Blue arrows represent joint axes, red lines represent muscle-tendon units (right half). (b) The model interacting with an exoskeleton in simulation. (c) The model with a prosthetic leg and a crutch.
Here, we propose a full-body human musculoskeletal model with 90 body segments, 206 joints and 700 muscle-tendon units for neuromuscular-driven dynamic simulation, as shown in Fig.1. The model can be imported into open-source physics simulation engines like MuJoCo[5] and OpenSim[6], establishing a framework that can be used to perform neuro-musculo-skeletal control simulation of human full-body movements.
@@
Our Contributions. We build a dynamic model of full-body human musculoskeletal system with physiologically feasible body segments, joints, and multi-muscle-tendon units. We embed the full-body model in simulation environments that support contact interaction simulations, providing a scalable, adaptable, and computationally efficient simulation framework for data-driven musculoskeletal system control. We propose a deep reinforcement learning algorithm, two-stage hierarchical training, for the control of the high-dimensional and highly nonlinear full-body musculoskeletal model. Our control algorithm can generate desired motions in full-body musculoskeletal model.
@@ II Related Work
Figure 2: Learning to walk with our full-body musculoskeletal model via control algorithm TSHT.
@@ IV Control
Figure 4: Overview of Two-Stage Hierarchical Training. In the collection stage, the agent determines muscle cluster actions. Muscle Synergy function then maps these muscle cluster actions into full-body muscle activation, leveraging prior physiological knowledge. Meanwhile, the reference generator generates the desired joint states based on motion capture data and computes corresponding rewards. Once a sufficient number of sub-optimal trajectories are collected, low-dimensional representations are extracted by encoder-decoder architecture and utilized for the training stage.
@@ V Experiments
We compare the TSHT algorithm on our full-body model with representative baselines, including SAR[35] and E2E RL. We design a series of full-body motion tasks and interaction scenarios to demonstrate the superiority of our model and algorithm.
@@ V-C Experimental Results
In our experiments, we utilized the SAC implementation from the open-source RL library stable-baselines3[47]. For the Walking forward and Human-Exo interaction tasks, both the policy network and the value function network took dimensions of (512, 300). For Prosthetic walking, the dimensions were (400, 300), while all other parameters remained at their default settings. All observation values were normalized to a maximum value of 10. For the Walking forward task, the collection stage has a length ($M$) of 15M, while the training stage ($N$) is 20M. In the Human-Exo interaction and Prosthetic walking tasks, both $M$ and $N$ are set to 15M. The latent dimension $d$ is 50.
@@ V-D Analysis and Discussion
Fig. 5(c) and Fig. 5(d) shows the representative joints’ angular position and muscle activity of the lower limbs in simulation. A walking simulation was presented along with the motion capture data and electromyography (EMG) recordings[12], which we consider as our reference. The periodicity and stability of the curves substantiate the robustness of our algorithm’s control efficacy. The discrepancy between the simulated and reference angular positions of the joints can be attributed to suboptimal reference trajectories that fail to ensure balance in the dynamics simulation of the full-body model. Consequently, an agent may opt for a more stable policy, such as increasing hip flexion and ankle dorsiflexion to elevate toe height and prevent tripping or falling, which is indicated in Fig. 5(c). The relaxation training phase after imitation[48] may be used to improve the dynamic motion performance. The temporal pattern of simulated muscle activations corresponds with key features of the EMG signals recorded during walking, despite exhibiting discrepancies in timing and magnitude for certain muscle activations when compared to EMG data. Achieving simulated activations that accurately replicate the measured EMG activity across all muscles remains a crucial challenge in studies utilizing muscle-driven simulations.[12, 17].
@@ VI Conclusion
This work introduces a musculoskeletal model including 90 rigid body segments, 206 joints, and 700 muscle-tendon units that encompasses the torso and limbs of human body. The model is adaptable to efficient open-source simulation engines, offering a simulation framework with a versatile interface for neuro-musculo-skeletal control simulation of human full-body movements. We introduce a novel control algorithm designed for high-dimensional and nonlinear musculoskeletal models to generate biologically plausible human motion. Our algorithm employs a hierarchical low-dimensional representation approach, integrating deep reinforcement learning and imitation techniques. Our proposed method exhibits superior performance in full-body musculoskeletal model control tasks, surpassing baseline algorithms significantly. Furthermore, the derived control policy of the agents provides valuable insights into the quantitative analysis of human motion control. Modeling and controlling musculoskeletal systems hold the potential to deepen our understanding of human motor intelligence and human factors in human-machine interactions. As a self-model of human for embodied intelligence, it could serve as a testing ground for the design of interactive robots and offer insights into humanoid behavior.

### v3

@@ Abstract
Modeling and control of the human musculoskeletal system is important for understanding human motor functions, developing embodied intelligence, and optimizing human-robot interaction systems. However, current human musculoskeletal models are restricted to a limited range of body parts and often with a reduced number of muscles. There is also a lack of algorithms capable of controlling over 600 muscles to generate reasonable human movements. To fill this gap, we build a musculoskeletal model (MS-Human-700) with 90 body segments, 206 joints, and 700 muscle-tendon units, allowing simulation of full-body dynamics and interaction with various devices. We develop a new algorithm using low-dimensional representation and hierarchical deep reinforcement learning to achieve state-of-the-art full-body control. We validate the effectiveness of our model and algorithm in simulations with real human locomotion data. The musculoskeletal model, along with its control algorithm, will be made available to the research community to promote a deeper understanding of human motion control and better design of interactive robots.
Project page: https://lnsgroup.cc/research/MS-Human-700
@@ I Introduction
Many of our daily movements require the use of the entire body. For instance, walking and running involve the coordinated movement of the lower limbs for propulsion, complemented by the swinging motion of the upper limbs for balance, while the trunk muscles contribute to torso stability. However, there is no musculoskeletal model that represents the entire human body. Constructing a full-body human musculoskeletal model enables comprehensive characterization of the action space and state space in human movement simulations. The dynamics of the model improves our understanding of human motor intelligence. Furthermore, current algorithms struggle to address the highly dimensional control problem associated with whole-body human musculoskeletal motion.
@@
Figure 1: (a) MS-Human-700: full-body musculoskeletal model. Blue arrows represent joint axes, red lines represent muscle-tendon units (right half). (b) The model interacting with an exoskeleton in simulation. (c) The model with a prosthetic leg and a crutch.
Here, we develop MS-Human-700, a full-body MusculoSkeletal Human model with 90 body segments, 206 joints, and 700 muscle-tendon units for neuromuscular-driven dynamic simulation, as shown in Fig.1. The model can be imported into open-source physics simulation engines like MuJoCo[5] and OpenSim[6], establishing a framework that can be used to perform neuro-musculo-skeletal control simulation of human full-body movements.
@@
Our Contributions. We build a dynamic model of full-body human musculoskeletal system with physiologically feasible body segments, joints, and multi-muscle-tendon units. We embed the whole-body model in simulation environments that support contact interaction simulations, providing a scalable, adaptable, and computationally efficient simulation framework for data-driven musculoskeletal system control. We propose a deep reinforcement learning algorithm, two-stage hierarchical training, for the control of the high-dimensional and highly nonlinear full-body musculoskeletal model. Our control algorithm can generate desired motions in full-body musculoskeletal model.
@@ II Related Work
Figure 2: Learning to walk with MS-Human-700 via Two-Stage Hierarchical Training (TSHT).
@@ IV Control
Figure 4: Two-Stage Hierarchical Training (TSHT). In the collection stage, the agent determines muscle cluster actions. Muscle Synergy function then maps these muscle cluster actions into full-body muscle activation, leveraging prior physiological knowledge. Meanwhile, the reference generator generates the desired joint states based on motion capture data and computes corresponding rewards. Once a sufficient number of sub-optimal trajectories are collected, low-dimensional representations are extracted by encoder-decoder architecture and utilized for the training stage.
@@ V Experiments
We compare the TSHT algorithm on MS-Human-700 model with representative baselines, including SAR[35] and E2E RL. We design a series of full-body motion tasks and interaction scenarios to demonstrate the superiority of our model and algorithm.
@@ V-C Experimental Results
In our experiments, we used the SAC implementation from the open-source RL library stable-baselines3[47]. For the Walking forward and Human-Exo interaction tasks, both the policy network and the value function network took dimensions of (512, 300). For Prosthetic walking, the dimensions were (400, 300), while all other parameters remained at their default settings. All observation values were normalized to a maximum value of 10. For the Walking forward task, the collection stage has a length ($M$) of 15M, while the training stage ($N$) is 20M. In the Human-Exo interaction and Prosthetic walking tasks, both $M$ and $N$ are set to 15M. The latent dimension $d$ is 50.
@@ V-D Analysis and Discussion
Fig. 6(c) and Fig. 6(d) shows the representative joints’ angular position and muscle activity of the lower limbs in simulation. A walking simulation was presented along with the motion capture data and electromyography (EMG) recordings[12], which we consider as our reference. The periodicity and stability of the curves substantiate the robustness of our algorithm’s control efficacy. The discrepancy between the simulated and reference angular positions of the joints can be attributed to suboptimal reference trajectories that fail to ensure balance in the dynamics simulation of the full-body model. Consequently, an agent may opt for a more stable policy, such as increasing hip flexion and ankle dorsiflexion to elevate toe height and prevent tripping or falling, which is indicated in Fig. 6(c). The relaxation training phase after imitation[48] may be used to improve the dynamic motion performance. The temporal pattern of simulated muscle activations corresponds with key features of the EMG signals recorded during walking, despite exhibiting discrepancies in timing and magnitude for certain muscle activations when compared to EMG data. Achieving simulated activations that accurately replicate the measured EMG activity across all muscles remains a crucial challenge in studies utilizing muscle-driven simulations[12, 17].
@@ VI Conclusion
This work presents a musculoskeletal model (MS-Human-700) with 90 rigid body segments, 206 joints, and 700 muscle-tendon units that encompasses the torso and limbs of human body. The model is adaptable to efficient open-source simulation engines, offering a simulation framework with a versatile interface for neuro-musculo-skeletal control simulation of human full-body movements. We introduce a novel control algorithm designed for high-dimensional and nonlinear musculoskeletal models to generate biologically plausible human motion. Our algorithm employs a hierarchical low-dimensional representation approach, integrating deep reinforcement learning and imitation techniques. Our proposed method exhibits superior performance in full-body musculoskeletal model control tasks, surpassing baseline algorithms significantly. Furthermore, the derived control policy of the agents provides valuable insights into the quantitative analysis of human motion control. Modeling and controlling musculoskeletal systems hold the potential to deepen our understanding of human motor intelligence and human factors in human-machine interactions. As a self-model of human for embodied intelligence, it could serve as a testing ground for the design of interactive robots and offer insights into humanoid behavior.